# The category of anyon sectors for non-abelian quantum double models

### v1

@@ 5.1 Finite ribbon multiplets
    where $B(-,-)$ is the braiding on $\mathbf{Rep}_{f}{\mathcal{D}}(G)$, and $P_{12}$ interchanges the factors in the tensor product of the representation spaces of $D_{1}$ and $D_{2}$ (see Appendix A).
@@ 5.2.1 Construction
For each site in the model, it is possible to define an action $\gamma:{\mathcal{D}}(G)\to\operatorname{Aut}(\mathcal{A})$ of the quantum double Hopf algebra. The amplimorphisms constructed here transform covariantly with respect to this action. These transformation properties (and of the ribbon multiplets themselves under this action) are essentially what connects these amplimorphisms to representations of ${\mathcal{D}}(G)$. For our purposes it is not necessary to spell out the details, and we refer the interested reader to [HamdanThesis].
@@ 5.2.3 Transportability
We say a half-infinite ribbon $\rho$ is ‘good’ if it is supported in a cone $\Lambda$ and for any other cone $\Lambda^{\prime}$ that is disjoint from $\Lambda$, there is a half-infinite ribbon $\rho^{\prime}$ and a bridge from $\rho$ to $\rho^{\prime}$. Note that any cone contains plenty of goof half-infinite ribbons.
@@
$(\pi_{0}\otimes\mathrm{id})\circ\mu_{\rho^{\prime}}^{D}=\mathrm{Ad}[U]\circ(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho}^{D}.$ (5.12)
@@
By Lemma C.2 the components of the image of $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ under $\pi_{0}\otimes\mathrm{id}$ converge in the strong-* topology, and therefore so does the full image of $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$. Denote the limit by $U$. Since the $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ are all unitary (Proposition 5.2) it follows from Lemma C.1 that $U$ is unitary. Since all the $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ are supported in the cone $\Lambda$, it follows that $U\in M_{n}({\mathcal{R}}(\Lambda))$.
@@
$U\cdot(\pi_{0}\otimes\mathrm{id})\big{(}\mu^{D}_{\rho}(O)\big{)}$ $=\lim_{n\uparrow\infty}\,(\pi_{0}\otimes\mathrm{id})\big{(}\mathbf{F}_{\bar{\sigma}_{n}}^{D}\cdot\mu^{D}_{\rho}(O)\big{)}$
$=\lim_{n\uparrow\infty}\,(\pi_{0}\otimes\mathrm{id})\big{(}\mu^{D}_{\rho^{(n)}}(O)\cdot\mathbf{F}_{\bar{\sigma}_{n}}^{D}\big{)}$
$=(\pi_{0}\otimes\mathrm{id})\big{(}\mu^{D}_{\rho^{\prime}}(O)\big{)}\cdot U$
where we used componentwise continuity of multiplication in the strong operator topology in the first equality, Eq. (5.13) to obtain the second equality, and the fact that $\mu^{D}_{\rho^{(n)}}(O)=\mu^{D}_{\rho^{\prime}}(O)$ for $n$ large enough and again componentwise continuity of multiplication to obtain the last equality. Since ${\mathcal{A}}^{\mathrm{loc}}$ is dense in ${\mathcal{A}}$. We conclude that Eq. (5.12) holds, which completes the proof. $\square$
@@
This answers a question that was left open in [Naaijkens2015], namely the construction of unitary charge transporters that transport charges between two cones, and not just over a finite distance.
@@ 5.3 Amplimorphisms of the allowed algebra from ribbon multiplets
$\chi_{\rho}^{D}\circ\pi_{0}(O)=(\pi_{0}\otimes\mathrm{id})\circ\mu_{\rho}^{D}(O).$ (5.14)
for all $O\in\mathcal{A}$. Moreover, $\chi_{\rho}^{D}$ is localised in $\Lambda$ and is transportable. It is therefore an object of $\mathbf{Amp}$.
Proof : First we extend $\mu_{\rho}^{D}$ to $\chi_{\rho}^{D}:{\mathcal{R}}(\Lambda)\rightarrow M_{n}({\mathcal{R}}(\Lambda))$. Note first that for every $\Lambda$ we have the existence of a forbidden cone $\widehat{\Lambda}$ disjoint from $\Lambda$. Since $\rho$ is good and by Lemma 5.8 we have $\mu_{\rho}^{D}\simeq\mu_{\hat{\rho}}^{D}$ where $\hat{\rho}$ is localised in $\widehat{\Lambda}$. Let $U$ be the unitary implementing this equivalence. Note that for all $O\in\mathcal{A}_{\Lambda}$ we have $\mu_{\hat{\rho}}^{D}(O)=O\otimes\mathds{1}_{n}$.
Define $\chi_{\rho}^{D}(O):=\mathrm{Ad}[U](O\otimes\mathds{1}_{n})$ for all $O\in{\mathcal{R}}(\Lambda)$. Since $\mathrm{Ad}[U]$ is weakly continuous, we have that $\chi_{\rho}^{D}$ is uniquely determined by $\mu_{\rho}^{D}$. This action on ${\mathcal{R}}(\Lambda)$ is independent of the choice of forbidden cone $\widehat{\Lambda}$, so the extensions to ${\mathcal{R}}(\widetilde{\Lambda})$ for different cones are consistent with each other. These actions therefore define a *-homomorphism $\chi_{\rho}^{D}$ on all of ${\mathcal{B}}$.
@@
Since this holds for any $O\in{\mathcal{A}}_{\Lambda^{c}}^{\mathrm{loc}}$, we find that $\chi_{\rho}^{D}$ is localised in $\Lambda$.
Now consider an allowed cone $\widetilde{\Lambda}$. Using transportability of $\mu_{\rho}^{D}$ we have some $\mu_{\tilde{\rho}}^{D}\simeq\mu_{\rho}^{D}$ localised in $\widetilde{\Lambda}$. Uniquely extend $\mu_{\tilde{\rho}}^{D}$ to $\chi_{\tilde{\rho}}^{D}$ as above. Then any unitary intertwiner from $\mu_{\tilde{\rho}}^{D}$ to $\mu_{\rho}^{D}$ is an equivalence between $\chi_{\rho}^{D}$ and $\chi^{D}_{\tilde{\rho}}$, showing that $\chi_{\rho}^{D}$ is indeed transportable. $\square$
@@ 6 Simple objects of $\mathbf{Amp}$
In order to do this we must first establish that the amplimorphisms $\chi_{\rho}^{D}$ are finite, so that they belong to $\mathbf{Amp}_{f}$. Then we must show that $\chi_{\rho}^{D}$ is a simple object whenever $D$ is an irreducible representation. Conversely, we must show that any simple object of $\mathbf{Amp}$ is equivalent to an amplimorphism $\chi^{D}_{\rho}$ for some irreducible representation $D$. In this section we prove these facts by appealing to the classification of irreducible anyon sectors of Kitaev’s quantum double models achieved in [bols2023classificationanyonsectorskitaevs], which we first review.
@@ 6.1 Classification of irreducible anyon sectors
for all $O\in{\mathcal{A}}_{\Lambda}$. If $\pi$ is moreover irreducible, then we call $\pi$ an anyon representation.
The following theorem follows directly from Theorem 2.4 and Proposition 5.19 of [bols2023classificationanyonsectorskitaevs].
Theorem 6.2 ([bols2023classificationanyonsectorskitaevs]).
@@ 7.1.1 Monoidal equivalence
where $I$ is the finite set of equivalence classes of irreducible representations of ${\mathcal{D}}$ and $D^{(r)}$ is a representation in class $r$. Let $u_{i}\in(D_{i}|\tilde{D}_{i})$ be the unitaries implementing these equivalences. It follows from Proposition 7.1 that $U_{i}=(\mathbbm{1}\otimes u_{i})\in(\chi_{\rho}^{D_{i}}|\chi_{\rho}^{\tilde{D}_{i}})$ and therefore $\tilde{T}:=U_{1}^{*}TU_{2}\in(\chi_{\rho}^{\tilde{D}_{1}}|\chi_{\rho}^{\tilde{D}_{2}})$.
@@
Let $\mathrm{id}_{D_{1},D_{2}}:F(D_{1})\otimes F(D_{2})\rightarrow F(D_{1}\times D_{2})$ be the identity maps. Strict monoidality of $F$ means that the $\mathrm{id}_{D_{1},D_{2}}$ form a natural transformation between functors ×$\times\circ(F,F):\mathbf{Rep}_{f}{\mathcal{D}}(G)\times\mathbf{Rep}_{f}{\mathcal{D}}(G)\rightarrow\mathbf{Amp}_{\rho}$ and $F\circ\times:\mathbf{Rep}_{f}{\mathcal{D}}(G)\times\mathbf{Rep}_{f}{\mathcal{D}}(G)\rightarrow\mathbf{Amp}_{\rho}$. Since $\mathbf{Amp}_{\rho}$ is strict, this boils down to $F(t)\times F(t^{\prime})=F(t\times t^{\prime})$ for any $t\in(D_{1}|D_{2})$ and any $t^{\prime}\in(D_{1}^{\prime}|D_{2}^{\prime})$, but this follows immediately from the definitions (recall in particular the definition (3.3) of the tensor product of intertwiners of amplimorphisms).
@@ 8 Conclusions
The problem is tractable for the quantum double model largely because the Hamiltonian is of commuting projector type. In general, we are interested in the whole quantum phase. The Hamiltonian of the quantum double model has a spectral gap in the thermodynamic limit, and roughly speaking another state is said to be in the same phase as the frustration free ground state $\omega_{0}$ of the quantum double model if they can be realised as ground states of a continuous path of gapped Hamiltonians. [^4] Alternatively, it is possible to give a definition of a phase without referring to Hamiltonians at all, using e.g. finite depth quantum circuits or suitable locality preserving automorphisms. Using standard techniques (which we outline below) our results carry over to other states in the same gapped phase, which may no longer be ground states of a commuting projector Hamiltonian. One of the features of the quantum double model is that the physical features should be stable against small perturbations. Indeed, the ground state has what is called local topological quantum order (LTQO) [Fiedler2014, Cui2020kitaevsquantum]. This implies that sufficiently small local perturbations (even if applied throughout the system) do not close the spectral gap [michalakisz, brahami:2010].
The result mentioned above implies that the ground states of the unperturbed and perturbed quantum double models can be related via an automorphism of $\mathcal{A}$ which is sufficiently local (meaning it satisfies a Lieb–Robinson type bound) [bachmannmns]. Hence one can consider the phase of a ground state as all states that can be connected via such a sufficiently local automorphism. It turns out that the braided category of anyon sectors is an invariant of such a phase (that is, each state in the phase supports the same type of anyons). This follows from the work of Ogata [Ogata2021] (see also [Ogata2021b] for a review), applied to the category $\mathbf{DHR}$ (or $\mathbf{DHR}_{f}$). Alternatively, one can apply the approximation techniques developed there (necessary because one is forced to replace Haag duality by a weaker, approximate version) directly to the amplimorphisms constructed here.

### v2

@@ 5.1 Finite ribbon multiplets
Throughout the rest of this manuscript the tensor product $\otimes$ of two matrices over $\mathcal{A}$} will always mean the usual matrix tensor product, while the tensor product $\otimes$ of an element of $\mathcal{A}$} with a matrix over $\mathbb{C}$} means the amplifying tensor product, yielding a matrix over $\mathcal{A}$}.
@@
    where B(,)B(-,-) is the braiding on $\mathbf{Rep}_{f}{\mathcal{D}}(G)$, and $P_{12}$ interchanges the factors in the tensor product of the representation spaces of $D_{1}$ and $D_{2}$ (see Appendix A).
@@ 5.2.1 Construction
For each site in the model, it is possible to define an action $\gamma:{\mathcal{D}}(G)\to\operatorname{Aut}(\mathcal{A})$ of the quantum double Hopf algebra. The amplimorphisms constructed here transform covariantly with respect to this action. These transformation properties (and of the ribbon multiplets themselves under this action) are essentially what connects these amplimorphisms to representations of ${\mathcal{D}}(G)$. For our purposes it is not necessary to spell out the details, and we refer the interested reader to [Ham24].
@@ 5.2.3 Transportability
We say a half-infinite ribbon $\rho$ is ‘good’ if it is supported in a cone $\Lambda$ and for any other cone $\Lambda^{\prime}$ that is disjoint from $\Lambda$, there is a half-infinite ribbon $\rho^{\prime}$ and a bridge from $\rho$ to $\rho^{\prime}$. Note that any cone contains plenty of good half-infinite ribbons, both positive and negative ones.
@@
$(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho^{\prime}}^{D}=\mathrm{Ad}[U]\circ(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho}^{D}.$ (5.12)
@@
By Lemma C.2 the components of the image of $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ under $\pi_{0}\otimes\mathrm{id}_{n}$ converge in the strong-* topology, and therefore so does the full image of $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$. Denote the limit by $U$. Since the $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ are all unitary (Proposition 5.2) it follows from Lemma C.1 that $U$ is unitary. Since all the $\mathbf{F}_{\bar{\sigma}_{n}}^{D}$ are supported in the cone $\Lambda$, it follows that $U\in M_{n}({\mathcal{R}}(\Lambda))$.
@@
$U\cdot(\pi_{0}\otimes\mathrm{id}_{n})\big(\mu^{D}_{\rho}(O)\big)$ $=\lim_{n\uparrow\infty}\,(\pi_{0}\otimes\mathrm{id}_{n})\big(\mathbf{F}_{\bar{\sigma}_{n}}^{D}\cdot\mu^{D}_{\rho}(O)\big)$
$=\lim_{n\uparrow\infty}\,(\pi_{0}\otimes\mathrm{id}_{n})\big(\mu^{D}_{\rho^{(n)}}(O)\cdot\mathbf{F}_{\bar{\sigma}_{n}}^{D}\big)$
$=(\pi_{0}\otimes\mathrm{id}_{n})\big(\mu^{D}_{\rho^{\prime}}(O)\big)\cdot U$
where we used componentwise continuity of multiplication in the strong operator topology in the first equality, Eq. (5.13) to obtain the second equality, and the fact that $\mu^{D}_{\rho^{(n)}}(O)=\mu^{D}_{\rho^{\prime}}(O)$ for $n$ large enough and again componentwise continuity of multiplication to obtain the last equality. Since ${\mathcal{A}}^{\mathrm{loc}}$ is dense in $\mathcal{A}$}, we conclude that Eq. (5.12) holds, which completes the proof. $\square$
@@
This answers a question that was left open in [Naa15], namely the construction of unitary charge transporters that transport charges between two cones, and not just over a finite distance. Note that Lemma 5.8 implies that the representation $(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho}^{D}$ satisfies a variant of the superselection criterion, where we have $(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho}^{D}\upharpoonright{\mathcal{A}}_{\Lambda^{c}}\cong n\cdot\pi_{0}\upharpoonright{\mathcal{A}}_{\Lambda^{c}}$. That is, instead of unitary equivalence as in (1.1), we have quasi-equivalence. As we shall see shortly, in the case at hand the two notions can be seen to coincide.
@@ 5.3 Amplimorphisms of the allowed algebra from ribbon multiplets
$\chi_{\rho}^{D}\circ\pi_{0}(O)=(\pi_{0}\otimes\mathrm{id}_{n})\circ\mu_{\rho}^{D}(O).$ (5.14)
for all $O\in\mathcal{A}$. Moreover, $\chi_{\rho}^{D}$ is localized in $\Lambda$ and is transportable. It is therefore an object of $\mathbf{Amp}$.
Proof : Recall that $\mathcal{B}$} is a direct limit of cone algebras ${\mathcal{R}}(\Lambda)$. Note that $\mu_{\rho}^{D}$ restricts to an amplimorphism $\mathcal{A}_{\Lambda}\to M_{n}(\mathcal{A}_{\Lambda})$. We show that we can extend this (on both sides) to ${\mathcal{R}}(\Lambda)$. This construction is compatible with the direct structure on the set of allowed cones, and hence defines an amplimorphism of $\mathcal{B}$}.
To see that we can extend $\mu_{\rho}^{D}$ (restricted to $\mathcal{A}_{\Lambda}$) to $\chi_{\rho}^{D}:{\mathcal{R}}(\Lambda)\rightarrow M_{n}({\mathcal{R}}(\Lambda))$, note first that for every $\Lambda$ we have the existence of a forbidden cone $\widehat{\Lambda}$ disjoint from $\Lambda$. Since $\rho$ is good and by Lemma 5.8, we have that $\mu_{\rho}^{D}\simeq\mu_{\hat{\rho}}^{D}$ where $\hat{\rho}$ is localized in $\widehat{\Lambda}$. Let $U$ be the unitary implementing this equivalence. By locality we have that for all $O\in\mathcal{A}_{\Lambda}$, it holds that $\mu_{\widehat{\rho}}^{D}(O)=O\otimes\mathds{1}_{n}$.
Define $\chi_{\rho}^{D}(O):=\mathrm{Ad}[U](O\otimes\mathds{1}_{n})$ for all $O\in{\mathcal{R}}(\Lambda)$. By construction, it follows that $\chi_{\rho}^{D}(O)=\mu_{\rho}^{D}(O)$ for all $O\in\mathcal{A}_{\Lambda}$. Let $O\in{\mathcal{R}}(\Lambda)$. Then there exist $\mathcal{A}_{\Lambda}\ni O_{\lambda}\to O$ weakly since $\mathcal{A}_{\Lambda}$ is weak-operator dense in ${\mathcal{R}}(\Lambda)$. Hence we have
$\lim_{\lambda}\mu_{\rho}^{D}(O_{\lambda})=\lim_{\lambda}\mathrm{Ad}[U](O_{\lambda}\otimes\mathds{1}_{n})=\mathrm{Ad}[U](O\otimes\mathds{1}_{n})=\chi_{\rho}^{D}(O),$
[…]
@@
Since this holds for any $O\in{\mathcal{A}}_{\Lambda^{c}}^{\mathrm{loc}}$, we find that $\chi_{\rho}^{D}$ is localized in $\Lambda$.
Now consider an allowed cone $\widetilde{\Lambda}$. Using transportability of $\mu_{\rho}^{D}$ (Lemma 5.8) we have that there exists some $\mu_{\widetilde{\rho}}^{D}\simeq\mu_{\rho}^{D}$ localized in $\widetilde{\Lambda}$. Uniquely extend $\mu_{\widetilde{\rho}}^{D}$ to $\chi_{\widetilde{\rho}}^{D}$ as above. Then any unitary intertwiner from $\mu_{\widetilde{\rho}}^{D}$ to $\mu_{\rho}^{D}$ is an equivalence between $\chi_{\rho}^{D}$ and $\chi^{D}_{\widetilde{\rho}}$, showing that $\chi_{\rho}^{D}$ is indeed transportable. $\square$
@@ 6 Simple objects of $\mathbf{Amp}$
In order to do this we must first establish that the amplimorphisms $\chi_{\rho}^{D}$ are finite, so that they belong to $\mathbf{Amp}_{f}$. Then we must show that $\chi_{\rho}^{D}$ is a simple object whenever $D$ is an irreducible representation. Conversely, we must show that any simple object of $\mathbf{Amp}$ is equivalent to an amplimorphism $\chi^{D}_{\rho}$ for some irreducible representation $D$. In this section we prove these facts by appealing to the classification of irreducible anyon sectors of Kitaev’s quantum double models achieved in [BV25], which we first review.
@@ 6.1 Classification of irreducible anyon sectors
for all $O\in{\mathcal{A}}_{\Lambda^{c}}$. If $\pi$ is moreover irreducible, then we call $\pi$ an anyon representation.
The following theorem follows directly from Theorem 2.4 and Proposition 5.19 of [BV25].
Theorem 6.2 ([BV25]).
@@ 7.1.1 Monoidal equivalence
where $I$ is the finite set of equivalence classes of irreducible representations of ${\mathcal{D}}(G)$ and $D^{(r)}$ is a representation in class $r$. Let $u_{i}\in(D_{i}|\tilde{D}_{i})$ be the unitaries implementing these equivalences. It follows from Proposition 7.1 that $U_{i}=(\mathbbm{1}\otimes u_{i})\in(\chi_{\rho}^{D_{i}}|\chi_{\rho}^{\tilde{D}_{i}})$ and therefore $\tilde{T}:=U_{1}^{*}TU_{2}\in(\chi_{\rho}^{\tilde{D}_{1}}|\chi_{\rho}^{\tilde{D}_{2}})$.
@@
Let $\mathrm{id}_{D_{1},D_{2}}:F(D_{1})\otimes F(D_{2})\rightarrow F(D_{1}\times D_{2})$ be the identity maps. Strict monoidality of $F$ means that the $\mathrm{id}_{D_{1},D_{2}}$ form a natural transformation between functors $\times\circ(F,F):\mathbf{Rep}_{f}{\mathcal{D}}(G)\times\mathbf{Rep}_{f}{\mathcal{D}}(G)\rightarrow\mathbf{Amp}_{\rho}$ and $F\circ\times:\mathbf{Rep}_{f}{\mathcal{D}}(G)\times\mathbf{Rep}_{f}{\mathcal{D}}(G)\rightarrow\mathbf{Amp}_{\rho}$. Since $\mathbf{Amp}_{\rho}$ is strict, this boils down to $F(t)\times F(t^{\prime})=F(t\times t^{\prime})$ for any $t\in(D_{1}|D_{2})$ and any $t^{\prime}\in(D_{1}^{\prime}|D_{2}^{\prime})$, but this follows immediately from the definitions (recall in particular the definition in equation (3.3) of the tensor product of intertwiners of amplimorphisms).
@@ 7.2 Equivalence of $\mathbf{Amp}_{\rho}$ and $\mathbf{Amp}_{f}$
Remark 7.11.
As mentioned previously, we restrict to the category $\mathbf{DHR}_{f}$. Since dualizable DHR endomorphisms are automatically finite (in our sense of the terminology) by [LR97], and all objects in the category $\mathbf{Amp}_{f}$ are dualizable, our results imply that the restriction of the category $\mathcal{O}_{\Lambda_{0}}$ (as defined by Ogata [Oga22]) to dualizable sectors (i.e., those who admit a conjugate) is precisely $\mathbf{Rep}_{f}{\mathcal{D}}(G)$. We do not expect that $\mathcal{O}_{\Lambda_{0}}$ has any objects which are not equivalent to (possibly infinite) direct sums of objects in $\mathbf{Amp}_{f}$. For example, any simple direct summand of any such an object would be equivalent to a simple object in $\mathbf{Amp}_{f}$.
@@ 8 Conclusions
The problem is tractable for the quantum double model largely because the Hamiltonian is of commuting projector type. In general, we are interested in the whole quantum phase. The Hamiltonian of the quantum double model has a spectral gap in the thermodynamic limit, and roughly speaking another state is said to be in the same phase as the frustration free ground state $\omega_{0}$ of the quantum double model if they can be realised as ground states of a continuous path of gapped Hamiltonians. [^7] Alternatively, it is possible to give a definition of a phase without referring to Hamiltonians at all, using e.g. finite depth quantum circuits or suitable locality preserving automorphisms. Using standard techniques (which we outline below) our results carry over to other states in the same gapped phase, which may no longer be ground states of a commuting projector Hamiltonian. One of the features of the quantum double model is that the physical features should be stable against small perturbations. Indeed, the ground state has what is called local topological quantum order (LTQO) [FN15, Cui+20]. This implies that sufficiently small local perturbations (even if applied throughout the system) do not close the spectral gap [MZ13, BHM10].
The result mentioned above implies that the ground states of the unperturbed and perturbed quantum double models can be related via an automorphism of $\mathcal{A}$ which is sufficiently local (meaning it satisfies a Lieb–Robinson type bound) [Bac+12]. Hence one can consider the phase of a ground state as all states that can be connected via such a sufficiently local automorphism. It turns out that the braided category of anyon sectors is an invariant of such a phase (that is, each state in the phase supports the same type of anyons). This follows from the work of Ogata [Oga22] (see also [Oga23] for a review), applied to the category $\mathbf{DHR}$ (or $\mathbf{DHR}_{f}$). Alternatively, one can apply the approximation techniques developed there (necessary because one is forced to replace Haag duality by a weaker, approximate version) directly to the amplimorphisms constructed here.